OpenStudy (anonymous):

Solve the following system of equations. 2x + 3y - z = 1 3x + y + 2z = 12 x + 2y - 3z = -5

OpenStudy (anonymous):

ick i would cheat

OpenStudy (anonymous):

otherwise it is a real pain to use elimination repeatedly do you have to do it step by step?

OpenStudy (anonymous):

2x+3y-z=0 ...(1) 3x+y+2z=12 ...(2) x+2y-3z=-5 ...(3) from (1) 2x+3y=z,substitute in (2) and (3) 3x+y+2(2x+3y)=12 or 7x+7y=12 ...(4) x+2y-3(2x+3y)=-5 -5x-7y=-5 ...(5) adding (4) and (5) 2x=7 x=7/2 from (4) 7*7/2+7y=12 49+14y=24 14y=24-49=-25 y=-25/14 substitute the values of x and y and get the value of z

OpenStudy (texaschic101):

2x + 3y - z = 1 3x + y + 2z = 12 -->(-3) ---------------- 2x + 3y - z = 1 -9x -3y - 6z = -36 (result of multiplying by -3) --------------- -7x - 7z = -35 3x + y + 2z = 12 -->(-2) x + 2y - 3z = -5 --------------- -6x - 2y - 4z = -24 (result of multiplying by -2) x + 2y - 3z = -5 ---------------add -5x - 7z = - 29 -7x - 7z = -35 -->(-1) -5x - 7z = -29 ---------------- 7x + 7z = 35 (result of multiplying by -1) -5x - 7z = -29 ----------------add 2x = 6 x = 3 7x + 7z = 35 7(3) + 7z = 35 21 + 7z = 35 7z = 35 - 21 7z = 14 z = 2 2x + 3y - z = 1 2(3) + 3y - 2 = 1 6 - 2 + 3y = 1 4 + 3y = 1 3y = 1 - 4 3y = -3 y = -1 check... 3x + y + 2z = 12 3(3) + (-1) + 2(2) = 12 9 - 1 + 4 = 12 8 + 4 = 12 12 = 12 (correct) SOLUTION : x = 3, y = -1, z = 2

OpenStudy (texaschic101):

@Brianna09201995 ...sorry it took so long

OpenStudy (texaschic101):

I kept messing up, but I finally got it :)

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

sorry i wrote the wrong statement